# Brill-Noether loci on an Enriques surface covered by a Jacobian Kummer surface

### v1

@@ 1. Introduction
In the last decades, classical Brill-Noether theory, concerning line bundles on projective curves, has been extended to higher rank bundles on varieties of arbitrary dimension. In [4], Costa and Miró-Roig introduced the Brill-Noether loci $W^{k}_{H}$, whose set theoretic support is formed by those bundles in $M_{H}$ having at least $k$ independent sections, where $M_{H}$ denotes the moduli space of $H$-stable vector bundles on the smooth polarized algebraic variety $(X,H)$ of some rank and some fixed Chern classes.
Standard questions of generalized Brill-Noether theory naturally arose from the ones of the classical theory and concern non-emptiness, irreducibility, dimension of the irreducible components or singularities of $W^{k}_{H}$. These problems are very far from being solved for an arbitrary variety. Little is known even in the surface case, despite the progress which has been made for the projective plane $\mathbb{P}^{2}$ [7], K3 surfaces [13, 14], Hirzebruch surfaces [5], ruled surfaces [3] or fibered surfaces [16]. We indicate [2] for a survey regarding higher rank Brill-Noether theory for surfaces.
In this note, we analyze the case of an Enriques surface $Y$ such that its K3 cover is a Jacobian Kummer surface and provide a way of choosing the Chern classes $c_{1},c_{2}$ such that first Brill-Noether locus $W^{1}_{H}(2;c_{1},c_{2})$ is non-empty and strictly contained in the moduli space $M_{Y,H}(2;c_{1},c_{2})$ for any polarization $H$. The reasons for considering the above setup are a good understanding of the Picard lattice of a Jacobian Kummer surface and a theorem of F. Takemoto as reinterpreted by H. Kim [12, Section 2] which gives a recipe to construct rank $2$ stable vector bundles on an Enriques surface by pushing forward line bundles on its K3 cover.
The outline of this note is as follows. We start by recalling in the first part of Section 2 the basics of higher rank Brill-Noether theory and then we discuss Jacobian Kummer surfaces. In Section 3 we present sufficient conditions for line bundles on a Jacobian Kummer surface in order to obtain a proper first Brill-Noether locus on the corresponding Enriques surface. This is the content of Theorem 3.3. After this, in Propositions 3.6 and 3.9, we give general examples of line bundles satisfying the hypothesis of Theorem 3.3. Some specific examples are also discussed.
Acknowledgements. The authors would like to thank professor Marian Aprodu for suggesting the problem and helpful discussions.
@@ 2.1. Generalized Brill-Noether theory
We denote by $M_{H}:=M_{S,H}(r;c_{1},c_{2})$ the moduli space of $H$-stable vector bundles on $S$ of rank $r$ with fixed Chern classes $c_{i}$, for $i=1,2$. We recall the following (see e.g. [9, Section 4.5]):
@@
[4, Theorem 2.3] Let $(S,H)$ and $M_{H}=M_{S,H}(r;c_{1},c_{2})$ as above. We assume the following further condition:
@@
Then, for any $k\geq 0$, there exists a determinantal variety $W_{H}^{k}(r;c_{1},c_{2})$, called the $k$-th Brill-Nother loci, such that
@@
$=$ $\dim(M_{H})-k(k-(r\chi\mathcal{(}{O}_{X})+\frac{1}{2}c_{1}\cdot K_{S}-\frac{1}{2}c_{1}^{2}+c_{2}))$
@@
As one might expect, the number $\rho_{H}^{k}(r;c_{1},c_{2})$ is called the generalized Brill-Noether number and it is the expected dimension of $W_{H}^{k}(r;c_{1}.c_{2})$.
@@
As noticed in [4, Corollary 2.3], the vanishing assumption ()($\ast$)( ∗ ) of Theorem 2.3 holds true when $(c_{1}.H)\geq(rK_{S}.H)$. The latter is true in particular if the first Chern class $c_{1}$ is effective and the canonical divisor on $S$ in numerically trivial. Thus, for the purpose of our note, we record the following:
@@
Let $Y$ be an Enriques surface (i.e. a smooth projective surface with $q(Y)=p_{a}(Y)=0$ and $02K_{Y}=$ 0) and $c_{i}\in H^{2i}(Y,\mathbb{Z})$, for $i=1,2$, such that $c_{1}$ is effective. Then, for any polarization $H$, the vanishing condition ()($\ast$)( ∗ ) holds true. Consequently, for any $r\geq 2$ and for any $k\geq 0$ there exists the $k$-th Brill-Noether locus $W^{k}_{H}(r;c_{1},c_{2})$.
@@ 2.2. Jacobian Kummer surfaces
In the sequel, we recall the construction of a Jacobian Kummer surface (see e.g. [10], [11]). Let $C$ be a smooth complex projective curve of genus $2$ and $\mathcal{A}=J(C)$ its Jacobian variety. Note that $\mathcal{A}$ is an abelian surface which comes with an involution $\iota:\mathcal{A}\rightarrow\mathcal{A}$ with $16$ fixed points. The quotient $\mathcal{K}=\mathcal{A}/\iota$ can be identified with the so-called Kummer quartic, which is a surface of degree $4$ in $\mathbb{P}^{3}$ with the maximal possible number of $16$ singular points. Each singular point is a node. Moreover, there are precisely $16$ planes in $\mathbb{P}^{3}$ touching $\mathcal{K}$ along a conic. Each of these conics is called a trope. An important feature of $\mathcal{K}$ is the so-called $(16_{6})$-configuration, that is any node lies on exactly $6$ tropes and each trope passes through exactly $6$ nodes.
The Jacobian Kummer surface $X=Kum(\mathcal{A})$ is the minimal desingularization of $\mathcal{K}$. The exceptional curves on $X$ lying over the nodes of $\mathcal{K}$ and the proper transforms of the tropes of $\mathcal{K}$ will be also called nodes and respectively tropes of $X$. The nodes and the tropes of the Jacobian Kummer surface $X$ form two families of $16$ mutually disjoint $(-2)$-curves on $X$. They can be described more explicitly in order to have a better understanding of the Picard lattice of $X$. For references we indicate [15, Section 4], [1, Section 2].
Take $p_{1},\ldots,p_{6}$ the Weierstrass points of $C$. The $2$-torsion points of $\mathcal{A}$ are $[0]$ and $[p_{i}-p_{j}]$, $1\leq i<j\leq 6$. Accordingly, we denote the nodes of $X$ by $E_{0}$ and $E_{ij}$, $1\leq i<j\leq 6$. The curve $C$ has also precisely $16$ theta caracteristics (i.e. divisor classes $D\in\operatorname{Pic}(C)$ such that $2D\sim K_{C}$), namely $[p_{i}]$, $1\leq i\leq 6$ and $[p_{i}+p_{j}-p_{6}]$ with $1\leq i<j\leq 5$, giving rise to $16$ theta divisors on $\mathcal{A}$, whose proper transforms are the tropes of $X$. Accordingly, we denote the tropes of $X$ by $T_{i}$, $1\leq i\leq 6$ and $T_{ij6}$, $1\leq i<j\leq 5$.
Let (.)(\ .\ )( . ) be the intersection form on $X$. The following lemma describes the Picard lattice of $X$ (cf. [15]):
@@
The Picard group of $X$ is generated over $\mathbb{Z}$ by $E_{0}$, $E_{ij}$, $T_{i}$, $T_{ij6}$ and
@@
Due to a result of Keum [10, Theorem 2], there exists a fixed-point-free involution $\theta$ of $X$, giving rise to an Enriques surface $Y=X/\theta$. Let us denote by $\pi:X\rightarrow Y$ the étale double cover of $Y$. The set of line bundles on $X$ which come from line bundles on $Y$ via pullback is described by a result of Horikawa [8, Theorem 5.1] and recasted in the following form in [1, Lemma 2.2]:
@@
The action of $\theta^{*}$ is presented in the table below [1]:
@@ 3.1. Main result
Let $X$ be a Jacobian Kummer surface, $\theta$ its fixed-point-free involution, $Y=X/\theta$ and $\pi:X\rightarrow Y$ the double cover of $Y$. In the sequel, we will state and prove the main result of this note. But first, let us recall the following important result of Takemoto, as recast by H. Kim [12, Section 2]:
@@
Secondly, let us prove that $W_{H}^{1}(2;c_{1},c_{2})\subsetneq M_{H}(2;c_{1},c_{2})$. Suppose by way of contradiction that $W_{H}^{1}(2;c_{1},c_{2})=M_{H}(2;c_{1},c_{2})$ and let $M_{0}$ be the irreducible component of the moduli space $M_{H}(2;c_{1},c_{2})$ containing $[\mathcal{V}]$. Since $h^{0}(\mathcal{V})=1$, there exists an open subset $U\subseteq M_{0}$ such that $h^{0}(\mathcal{E})=1$ for any $[\mathcal{E}]\in U$. Note that for any $[\mathcal{E}]\in U$, the vanishing locus $V(s)$ of the unique nonzero section $s\in H^{0}(\mathcal{E})$ has a divisorial part $C_{\mathcal{E}}$, which may be zero, and a zero-dimensional part $Z_{\mathcal{E}}$. Since the Picard group of $Y$ is discrete, we may take an open subset $U^{\prime}\subseteq U$ such that $C_{\mathcal{E}}=C$ for any $[\mathcal{E}]\in U^{\prime}$.
@@
Now let $\mathcal{P}$ be the projective bundle over $\operatorname{Hilb}_{c_{2}}(Y)$, whose fiber over the point $Z\in\operatorname{Hilb}_{c_{2}}(Y)$ is the space of weak isomorphism classes of extensions of type:
@@
Claim 2: $\dim\mathcal{P}\leq 3c_{2}-c_{1}^{2}/2-1$.
Proof of Claim 2: Let us take a generic point $[\mathcal{E}]$ in $U^{\prime}$ and $Z\in\operatorname{Hilb}_{c_{2}}(Y)$ the vanishing locus of the unique section of $\mathcal{E}$. Let us also denote $\mathcal{J}=\mathcal{I}_{Z}(c_{1})\otimes K_{Y}$. After twisting the exact sequence
$0\rightarrow\mathcal{I}_{Z}\rightarrow\mathcal{O}_{X}\rightarrow\mathcal{O}_{Z}\rightarrow 0$
@@
On the other hand, from the exact sequence
@@
we deduce that $h^{0}(\mathcal{J})=h^{0}(\mathcal{E}\otimes K_{Y})$. But since $\mathcal{E}$ is generic
$h^{0}(\mathcal{E}\otimes K_{Y})\leq h^{0}(\pi_{\ast}\mathcal{O}_{X}(D)\otimes K_{Y})$
and then
$\begin{split}h^{0}(\pi_{\ast}\mathcal{O}_{X}(D)\otimes K_{Y})&=h^{0}(\pi_{\ast}(\mathcal{O}_{X}(D)\otimes\pi^{\ast}K_{Y}))\\ &=h^{0}(\mathcal{O}_{X}(D)\otimes\pi^{\ast}K_{Y})\\ &=h^{0}(\mathcal{O}_{X}(D)\otimes\mathcal{O}_{X})=1\end{split}$
We also note that $h^{2}(\mathcal{E}\otimes K_{Y})=h^{0}(\mathcal{E^{\vee}})=0$ and thus, using the last exact sequence, we obtain the vanishing of $h^{2}(\mathcal{J})$. Therefore, formula (2) ensures us that:
$h^{1}(\mathcal{J})\leq c_{2}-\frac{c_{1}^{2}}{2}$
@@
Finally, by [6, Propositions 27, 28] one has that $c_{1}^{2}=[\pi_{*}D]^{2}$ and $c_{2}=\frac{1}{2}([\pi_{*}D]^{2}-\pi_{*}(D^{2}))=\frac{1}{2}[\pi_{*}D]^{2}-\frac{1}{2}(D^{2})$. Hence we obtain that $c_{2}-\frac{1}{2}c_{1}^{2}=-\frac{1}{2}D^{2}>2$, which is a contradiction. ∎
@@
Let $D$ be a nonzero effective divisor on $X$ and let $\mathcal{V}=\pi_{*}\mathcal{O}_{X}(D)$ and $c_{i}=c_{i}(\mathcal{V})$ for $i=1,2$. Then $h^{0}(\mathcal{E})\geq(D^{2})/2+2$, for any $[\mathcal{E}]\in M_{H}(2;c_{1},c_{2})$. Indeed, by Remark 2.4, $h^{2}(\mathcal{E})=0$, hence $h^{0}(\mathcal{E})\geq\chi(\mathcal{E})=\chi(\mathcal{V})=(D^{2})/2+2$. Therefore, if $D$ is either a node or a trope, then $(D)^{2}=-2$ and so $h^{0}(\mathcal{E})\geq 1$ for any $[\mathcal{E}]\in M_{H}(2;c_{1},c_{2})$, which shows that $W_{H}^{1}(2;c_{1},c_{2})=M_{H}(2;c_{1},c_{2})$ in this case.

### v2

@@ 1. Introduction
In the last decades, classical Brill-Noether theory, concerning line bundles on projective curves, has been extended to higher rank bundles on varieties of arbitrary dimension. In [CMR10a], Costa and Miró-Roig introduced the Brill-Noether loci $W^{k}_{H}$, whose set theoretic support is formed by those bundles in $M_{H}$ having at least $k$ independent sections, where $M_{H}$ denotes the moduli space of $H$-stable vector bundles of some rank and some fixed Chern classes on a smooth polarized algebraic variety $(X,H)$. In a recent paper, Nugent [Nug] provided a new construction of the Brill-Noether loci, generalizing the one in [CMR10a].
Standard questions of generalized Brill-Noether theory naturally arose from the ones of the classical theory and concern non-emptiness, irreducibility, dimension of the irreducible components or singularities of $W^{k}_{H}$. These problems are very far from being solved for an arbitrary variety. Little is known even in the surface case, despite the progress which has been made for the projective plane $\mathbb{P}^{2}$ [GLL23], [CHR], K3 surfaces [Ley12, Ley], Hirzebruch surfaces [CMR10b], ruled surfaces [CMT24] or fibered surfaces [RRT22]. We also indicate [CHN24] for a survey regarding higher rank Brill-Noether theory on surfaces.
In this note, we analyze the case of an Enriques surface $Y$ such that its K3 cover is a Jacobian Kummer surface and provide a way of choosing the Chern classes $c_{1},c_{2}$ such that first Brill-Noether locus $W^{1}_{H}(2;c_{1},c_{2})$ is non-empty and strictly contained in the moduli space $M_{Y,H}(2;c_{1},c_{2})$ for any polarization $H$. The reasons for considering the above setup are a good understanding of the Picard lattice of a Jacobian Kummer surface and a theorem of F. Takemoto as reinterpreted by H. Kim [Kim98, Section 2] which gives a recipe to construct rank $2$ stable vector bundles on an Enriques surface by pushing forward line bundles on its K3 cover.
The outline of this note is as follows. We start by recalling in the first part of Section 2 the basics of higher rank Brill-Noether theory and then we discuss Jacobian Kummer surfaces. In Section 3 we present sufficient conditions for line bundles on a Jacobian Kummer surface in order to obtain a proper first Brill-Noether locus on the corresponding Enriques surface. This is the content of Theorem 3.3. After this, in Propositions 3.7 and 3.15, we give general examples of line bundles satisfying the hypothesis of Theorem 3.3. Some specific examples are also discussed.
Acknowledgements. The authors would like to thank professor Marian Aprodu for suggesting the problem and helpful discussions and the referee for carefully reading the paper and valuable comments.
@@ 2.1. Generalized Brill-Noether theory
We denote by $M_{H}:=M_{S,H}(r;c_{1},c_{2})$ the moduli space of $H$-stable vector bundles on $S$ of rank $r$ with fixed Chern classes $c_{i}$, for $i=1,2$. We recall the following (see e.g. [HL10, Section 4.5]):
@@
[CMR10a, Theorem 2.3] Let $(S,H)$ and $M_{H}=M_{S,H}(r;c_{1},c_{2})$ as above. We assume the following further condition:
@@
Then, for any $k\geq 0$, there exists a determinantal variety $W_{H}^{k}(r;c_{1},c_{2})$, called the $k$-th Brill-Noether loci, such that
@@
$=$ $\dim(M_{H})-k(k-(r\chi\mathcal{(}{O}_{S})-\frac{1}{2}c_{1}\cdot K_{S}+\frac{1}{2}c_{1}^{2}-c_{2}))$
@@
As one might expect, the number $\rho_{H}^{k}(r;c_{1},c_{2})$ is called the generalized Brill-Noether number or the expected dimension of $W_{H}^{k}(r;c_{1}.c_{2})$.
@@
As noticed in [CMR10a, Corollary 2.3], the vanishing assumption ()($\ast$)( ∗ ) of Theorem 2.3 holds true when $(c_{1}.H)\geq(rK_{S}.H)$. The latter is true in particular if the first Chern class $c_{1}$ is effective and the canonical divisor on $S$ in numerically trivial. Thus, for the purpose of our note, we record the following:
@@
Let $Y$ be an Enriques surface (i.e. a smooth projective surface with $q(Y)=p_{a}(Y)=0$ and $02K_{Y}=$ 0) and $c_{i}\in H^{2i}(Y,\mathbb{Z})$, for $i=1,2$, such that $(c_{1}.H)>0$ for any polarization $H$. Then, for any polarization $H$, the vanishing condition ()($\ast$)( ∗ ) holds true because otherwise there would be an injective morphism $\mathcal{O}_{Y}\rightarrow\mathcal{E}^{\vee}\otimes K_{Y}$ and from here the vanishing can be proved quickly using stability and Chern classes. A particular case which will be of interest for us in this paper is when $c_{1}$ is effective, then $(c_{1}.H)>0$ for any polarization. Consequently, for any $r\geq 2$ and for any $k\geq 0$, the $k$-th Brill-Noether locus $W^{k}_{H}(r;c_{1},c_{2})$ exists.
@@ 2.2. Jacobian Kummer surfaces
In the sequel, we recall the construction of a Jacobian Kummer surface (see e.g. [Keu97, Section 1]). Let $C$ be a smooth complex projective curve of genus $2$ and $\mathcal{A}=J(C)$ its Jacobian variety. Note that $\mathcal{A}$ is an abelian surface which comes with an involution $\iota:\mathcal{A}\rightarrow\mathcal{A}$ with $16$ fixed points. The quotient $\mathcal{K}=\mathcal{A}/\iota$ can be identified with the so-called Kummer quartic, which is a surface of degree $4$ in $\mathbb{P}^{3}$ with the maximal possible number of $16$ singular points. Each singular point is a node. Moreover, there are precisely $16$ planes in $\mathbb{P}^{3}$ touching $\mathcal{K}$ along a conic. Each of these conics is called a trope. An important feature of $\mathcal{K}$ is the so-called $(16_{6})$-configuration, that is any node lies on exactly $6$ tropes and each trope passes through exactly $6$ nodes.
The Jacobian Kummer surface $X=Kum(\mathcal{A})$ is the minimal desingularization of $\mathcal{K}$. The exceptional curves on $X$ lying over the nodes of $\mathcal{K}$ and the proper transforms of the tropes of $\mathcal{K}$ will be also called nodes and respectively tropes of $X$. The nodes and the tropes of the Jacobian Kummer surface $X$ form two families of $16$ mutually disjoint $(-2)$-curves on $X$. They can be described more explicitly in order to have a better understanding of the Picard lattice of $X$. For references we indicate [Oha09, Section 4], [AK20, Section 2].
Take $p_{1},\ldots,p_{6}$ the Weierstrass points of $C$. The $2$-torsion points of $\mathcal{A}$ are $[0]$ and $[p_{i}-p_{j}]$, $1\leq i<j\leq 6$. Accordingly, we denote the nodes of $X$ by $E_{0}$ and $E_{ij}$, $1\leq i<j\leq 6$. The curve $C$ has also precisely $16$ theta characteristics (i.e. divisor classes $D\in\operatorname{Pic}(C)$ such that $2D\sim K_{C}$), namely $[p_{i}]$, $1\leq i\leq 6$ and $[p_{i}+p_{j}-p_{6}]$ with $1\leq i<j\leq 5$, giving rise to $16$ theta divisors on $\mathcal{A}$, whose proper transforms are the tropes of $X$. Accordingly, we denote the tropes of $X$ by $T_{i}$, $1\leq i\leq 6$ and $T_{ij6}$, $1\leq i<j\leq 5$.
The lemma below can be easily deduced from [Keu97, Section 1] and [AK20, Section 2]:
@@
The intersection form (.)(\ .\ )( . ) on $X$ has the following values for the nodes and the tropes of $X$:
  • 1)
@@
Due to a result of Keum [Keu90, Theorem 2], there exists a fixed-point-free involution $\theta$ of $X$, giving rise to an Enriques surface $Y=X/\theta$. Let us denote by $\pi:X\rightarrow Y$ the étale double cover of $Y$. The set of line bundles on $X$ which come from line bundles on $Y$ via pullback is described by a result of Horikawa [Hor78, Theorem 5.1] and recast in the following form in [AK20, Lemma 2.2]:
@@
The action of $\theta^{*}$ is presented in the table below [AK20]:
@@ 3.1. Main result
Let $X$ be a Jacobian Kummer surface, $\theta$ its fixed-point-free involution, $Y=X/\theta$ and $\pi:X\rightarrow Y$ the double cover of $Y$. In the sequel, we will state and prove the main result of this note. But first, let us recall the following important result of Takemoto, as recast by H. Kim [Kim98, Section 2]:
@@
Secondly, let us prove that $W_{H}^{1}(2;c_{1},c_{2})\subsetneq M_{H}(2;c_{1},c_{2})$. Suppose by way of contradiction that $W_{H}^{1}(2;c_{1},c_{2})=M_{H}(2;c_{1},c_{2})$ and let $M_{0}$ be the irreducible component of the moduli space $M_{H}(2;c_{1},c_{2})$ containing $[\mathcal{V}]$. Since $h^{0}(\mathcal{V})=1$, there exists an open subset $U\subseteq M_{0}$ such that $h^{0}(\mathcal{E})=1$ for any $[\mathcal{E}]\in U$. Note that for any $[\mathcal{E}]\in U$, the vanishing locus $V(s)$ of the unique nonzero section $s\in H^{0}(\mathcal{E})$ has a divisorial part $C_{\mathcal{E}}$, which may be zero, and a zero-dimensional part $Z_{\mathcal{E}}$. Since the Picard group of $Y$ is discrete and $h^{0}(C_{\mathcal{E}})=1$ for any $[\mathcal{E}]\in U$, we may take an open subset $U^{\prime}\subseteq U$ such that $C_{\mathcal{E}}=C$ for any $[\mathcal{E}]\in U^{\prime}$.
@@
Now let $\mathcal{P}$ be the projective bundle over $\operatorname{Hilb}_{c_{2}}(Y)$, whose fiber over the point $Z\in\operatorname{Hilb}_{c_{2}}(Y)$ is the space of weak isomorphism classes (see [Fri98, p. 31]) of extensions of type:
@@
Claim 2: $\dim\mathcal{P}=3c_{2}-c_{1}^{2}/2-2$.
Proof of Claim 2: Let us take a general point $[\mathcal{E}]$ in $U^{\prime}$ and $Z\in\operatorname{Hilb}_{c_{2}}(Y)$ the vanishing locus of the unique section of $\mathcal{E}$. Let us also denote $\mathcal{J}=\mathcal{I}_{Z}(c_{1})\otimes K_{Y}$. After twisting the exact sequence
$0\rightarrow\mathcal{I}_{Z}\rightarrow\mathcal{O}_{Y}\rightarrow\mathcal{O}_{Z}\rightarrow 0$
@@
Note that $h^{0}(\mathcal{J})=0$. Otherwise, denoting by $D_{Y}$ an effective divisor representing $c_{1}+K_{Y}$, we would have an injective map $\mathcal{I}_{D_{Y}}=\mathcal{O}_{Y}(-D_{Y})\hookrightarrow\mathcal{I}_{Z}$, and thus an inclusion $\operatorname{Supp}(Z)\subseteq\operatorname{Supp}(D_{Y})$, which is false, since $\mathcal{E}$ is general, hence so is $Z$ in $\operatorname{Hilb}_{c_{2}}(Y)$.
We also notice that, since $\mathcal{E}^{\vee}$ is $H$-stable of negative slope, by Serre duality we get $h^{2}(\mathcal{E}\otimes K_{Y})=h^{0}(\mathcal{E^{\vee}})=0$. Thus, using the exact sequence
@@
we obtain the vanishing of $h^{2}(\mathcal{J})$. Therefore, formula (2) ensures us that:
$h^{1}(\mathcal{J})=c_{2}-\frac{c_{1}^{2}}{2}-1$
@@
Finally, by [Fri98, Propositions 27, 28] one has that $c_{1}^{2}=[\pi_{*}D]^{2}$ and $c_{2}=\frac{1}{2}([\pi_{*}D]^{2}-\pi_{*}(D^{2}))=\frac{1}{2}[\pi_{*}D]^{2}-\frac{1}{2}(D^{2})$. Hence we obtain that $c_{2}-\frac{1}{2}c_{1}^{2}=-\frac{1}{2}(D^{2})\geq 2$, which is a contradiction. ∎
@@
Remark 3.5.
It is worth mentioning that the expected dimension of the Brill-Noether locus constructed in Theorem 3.3 is $\rho^{1}_{H}(2;c_{1},c_{2})=\operatorname{dim}M_{H}+1+\frac{1}{2}(D^{2})$.
@@
Let $D$ be a nonzero effective divisor on $X$ and let $\mathcal{V}=\pi_{*}\mathcal{O}_{X}(D)$ and $c_{i}=c_{i}(\mathcal{V})$ for $i=1,2$. Then $h^{0}(\mathcal{E})\geq(D^{2})/2+2$, for any $[\mathcal{E}]\in M_{H}(2;c_{1},c_{2})$. Indeed, by Remark 2.4, $h^{2}(\mathcal{E})=0$, hence $h^{0}(\mathcal{E})\geq\chi(\mathcal{E})=\chi(\mathcal{V})=(D^{2})/2+2$. Therefore, if $D$ is either a node or a trope, then $(D^{2})=-2$ and so $h^{0}(\mathcal{E})\geq 1$ for any $[\mathcal{E}]\in M_{H}(2;c_{1},c_{2})$, which shows that $W_{H}^{1}(2;c_{1},c_{2})=M_{H}(2;c_{1},c_{2})$ in this case. This fact can also be seen in the virtue of Theorem 2.3 and Remark 3.5.